I cleaned up around the 40 m lab. All the Laser Safety Glasses have been picked up and placed on the rack at the entrance.
Some miscellaneous BNC Connector cables have been arranged and organized along the wall parallel to the Y-Tunnel.
Nitrogen tanks have been swapped out. Current tank is at 1200 psi and the other is at 1850 psi.
The tool box has been organized with each tool in its specified area.